# Certifying rings of integers in number fields

### v1

@@ 4.6. The non-monogenic case
In this section, we specialize to $R=\mathbf{Z}$, our main interest. Our formalization efforts were primarily focused on this case, with $\mathbf{F}_{p}$ modeled as ZMod p. However, we formalized many theorems about semilinear maps (introduced in Lean in (Dupuis et al., 2022)) and their kernels –which we use extensively– in a more general setting, laying the groundwork for future extensions.
The strategy used in the Round-2 algorithm is to reduce the computation of $I_{p}$ and $r(I_{p})$ to a computation in linear algebra over the field $\mathbf{F}_{p}$. We will use an analogous strategy to certify $p$-maximality by verifying that the kernel of certain linear map is trivial. Let $\mathcal{O}$ be a commutative ring which is free as a $\mathbf{Z}$-module and of finite rank $r$, and let $p$ be a prime number. The quotient ring $\mathcal{O}/p\mathcal{O}$ is a finitely dimensional $\mathbf{F}_{p}$-algebra. Similarly, the quotient $I_{p}/pI_{p}$ is a finitely dimensional $\mathbf{F}_{p}$-module (albeit not an $\mathbf{F}_{p}$- algebra). Consider the map from $\mathcal{O}/p\mathcal{O}$ to the $\mathbf{F}_{p}$-linear endomorphisms of $I_{p}/pI_{p}$.
@@
Let $K$ be $\mathbf{Z}$-torsion free ring containing $\mathcal{O}$. If the kernel of the map $\varphi$ in (3) is trivial, then $r(I_{p})=\mathcal{O}$.
When $K$ is a $\mathbf{Q}$-algebra, this is an if-and-only-if statement. If $\mathcal{O}_{m}$ is a subring of $K$ containing $\mathcal{O}$ and of equal rank, then the previous proposition together with Theorem (4.1) ensures that $\mathcal{O}$ is $p$-maximal if the kernel of $\varphi$ is trivial. Therefore, finding a way to certify that $\varphi$ has trivial kernel becomes our main goal.
@@
Thus, $I_{p}/pI_{p}$ (taking $M=I_{p}$ and $R=\mathbf{Z}$) and $O/pO$ (taking $M=\mathcal{O}$ and $R=\mathcal{O}$) are both special cases of this construction, allowing Lean to infer an $\mathbf{F}_{p}$-module instance for each. Furthermore, Lean will automatically synthesis a CommRing instance for $\mathcal{O}/p\mathcal{O}$, which we then use to give it a $\mathbf{F}_{p}$-algebra structure. Another advantage of this unified approach is that we can define associated objects –such as a $\mathbf{Z}/n\mathbf{Z}$-basis for $M/nM$ given a $\mathbf{Z}$-basis for $M$, and apply it to both cases.
To define map (3) in Lean, we used multiple steps: First, for a parameter : $\mathcal{O}$), we defined the map $I_{p}\to I_{p}/pI_{p}$ sending $\beta\mapsto\overline{\alpha\beta}$. Using Quotient.lift we lifted this to a map $I_{p}/pI_{p}\to I_{p}/pI_{p}$. This was bundled with its proofs of linearity into a linear function, giving an element in $\operatorname{End}(I_{p}/pI_{p})$. After defining the corresponding map $\mathcal{O}\to\operatorname{End}(I_{p}/pI_{p})$, lifting again, and proving linearity, we obtain:
@@
Here, the typeclass system struggles finding the $\mathbf{F}_{p}$-module instance for the space of $\mathbf{F}_{p}$-linear maps $I_{p}/pI_{p}\rightarrow I_{p}/pI_{p}$. One can either increase the instance synthesis heartbeats limit or exclude the problematic path from the search, which is what we did.
Let us return to the matter of certifying that the kernel of $\varphi$ is trivial. Given a $\mathbf{Z}$-basis for an abelian additive group $M$, we obtain a $\mathbf{Z}/n\mathbf{Z}$-basis for $M/nM$ by mapping it through the $\sigma$-semilinear map $M\to M/nM$, with $\sigma:\mathbf{Z}\to\mathbf{Z}/n\mathbf{Z}$. With these bases, the image of an element in $M$ represented by $(x_{1},\ldots,x_{r})\in\mathbf{Z}^{r}$ will have coordinates $(\bar{x}_{1},\ldots,\bar{x}_{r})\in(\mathbf{Z}/n\mathbf{Z})^{r}$ in $M/nM$. In this way, we obtain an $\mathbf{F}_{p}$-basis for $\mathcal{O}/p\mathcal{O}$ from a $\mathbf{Z}$-basis for $\mathcal{O}$. Finding an $\mathbf{F}_{p}$-basis for $I_{p}/pI_{p}$ is more complex. We will make use of the Frobenius endomorphism $\operatorname{Frob}_{p}$, which, in a ring of characteristic $p$, is the $\mathbf{F}_{p}$-linear map $x\mapsto x^{p}$. For $j\in\mathbf{N}$, the iterated Frobenius endomorphism $\operatorname{Frob}_{p}^{j}$ sends $x\mapsto x^{p^{j}}$. The image of $I_{p}$ under the quotient map $\mathcal{O}\to\mathcal{O}/p\mathcal{O}$ is $I_{p}/p\mathcal{O}$, an $\mathbf{F}_{p}$-subspace of $\mathcal{O}/p\mathcal{O}$. The following proposition lets us realize $I_{p}/p\mathcal{O}$ as the kernel of a linear map:
@@
Let $j$ be an integer such $r\leq p^{j}$. The kernel of $\operatorname{Frob}_{p}^{j}:\mathcal{O}/p\mathcal{O}\to\mathcal{O}/p\mathcal{O}$ is equal to $I_{p}/p\mathcal{O}$.
While we could theoretically find a matrix representation for this map, an algorithm for computing the kernel of a matrix has not yet, as far as we know, been implemented in Lean. However, we can certify that a given set of elements in $\mathcal{O}/p\mathcal{O}$ is a basis for the kernel of $\operatorname{Frob}_{p}^{j}$ as follows:
Denote by $\bar{x}$ the image in $\mathcal{O}/p\mathcal{O}$ of $x$ in $\mathcal{O}$. Provide multisets $\mathcal{V}=\{v_{1},\ldots,v_{m}\}$ and $\mathcal{W}=\{w_{1},\ldots,w_{n}\}$ with elements in $\mathcal{O}$ –where $r=m+n$– such that $\overline{\mathcal{V}}=\{\bar{v}_{1},\ldots,\bar{v}_{m}\}$ and $\mathcal{U}=\{\operatorname{Frob}_{p}^{j}(\bar{w}_{1}),\ldots,\operatorname{Frob}_{p}^{j}(\bar{w}_{n})\}$ are $\mathbf{F}_{p}$-linearly independent and $\operatorname{Frob}_{p}^{j}(\bar{v}_{i})=0$ for all $i$. By a dimension argument, this guarantees that $\{\bar{v}_{1},\ldots,\bar{v}_{m}\}$ is an $\mathbf{F}_{p}$-basis for the kernel of $\operatorname{Frob}_{p}^{j}$ – and thus for $I_{p}/p\mathcal{O}$
We can represent $\overline{\mathcal{V}}$ as an $m\times r$ matrix over $\mathbf{F}_{p}$, the rows being the coordinates of the elements in $\overline{\mathcal{V}}$ with respect to the basis for $\mathcal{O}/p\mathcal{O}$. Since we are working over a field, we can choose the elements in $\mathcal{V}$ so that the corresponding matrix is in reduced row echelon form. The linear independence of $\mathcal{V}$ can then be read off directly, without any determinant calculation (which is rather slow using the current Lean implementation). Similarly, the right choice of $\mathcal{W}$ will make the linear independence of $\mathcal{U}$ immediate to determine.
@@
With a basis for $I_{p}/pI_{p}$, we can now represent an element in this space as an $r$-tuple, and an element in $\operatorname{End}(I_{p}/p_{p})$ as an $r\times r$ matrix over $\mathbf{F}_{p}$. To certify that the kernel of the map $\varphi$ in (3) is trivial, we can provide a multiset $\mathcal{X}=\{x_{1},\ldots,x_{r}\}$ of elements in $\mathcal{O}$ such that $\varphi(\overline{\mathcal{X}})=\{\varphi(\bar{x_{1}}),\ldots,\varphi(\bar{x_{m}}))\}$ is linearly independent. This last verification can be simplified by observing that, generically, we expect there to be an element in $\beta_{W}\in I_{p}/pI_{p}$ such that $\{\varphi(\bar{x}_{1})(\beta_{W}),\ldots,\varphi(\bar{x}_{m})(\beta_{W}))\}$ is linearly independent. This $\beta_{W}$ acts as a witness of the linear independence of $\varphi(\overline{\mathcal{X}})$. Furthermore, we can choose $\mathcal{X}$ in a way that makes the linear independence of the multiset easy to verify. This leads us to the following certificate for $p$-maximality.
@@
Consider $\mathcal{O}$ and $\mathcal{O}_{m}$ subalgebras of a $\mathbf{Z}$-torsion free commutative ring such that $\mathcal{O}\subseteq\mathcal{O}_{m}$, and both of rank $r$. Let =$\mathcal{B}=\{b_{1},\ldots,b_{r}\}$ be a $\mathbf{Z}$-basis for $\mathcal{O}$. Let $m,n$ and $t$ be non-negative integers with $r=m+n$ and $r\leq p^{t}$. A certificate for $p$-maximality consists of the following data: $m\times r$ matrix $\mathcal{V}$ over $\mathbf{Z}$ $n\times r$ matrix $\mathcal{W}$ over $\mathbf{Z}$ $n\times r$ matrix $\mathcal{U}$ over $\mathbf{F}_{p}$ $m$-tuple $\nu$ over $\mathbf{Z}_{>0}$ $n$-tuple $\omega$ over $\mathbf{Z}_{>0}$ $r\times r$ matrix $\mathcal{X}$ over $\mathbf{Z}$ $m$-tuple $\beta$ over $\mathbf{Z}$ $n$-tuple $\gamma$ over $\mathbf{Z}$ $r\times m$ matrix $\mathbf{a}$ over $\mathbf{Z}$. $r\times n$ matrix $\mathbf{c}$ over $\mathbf{Z}$.
@@
    • If $\eta_{i}$ is in $A$: $\bar{a}_{i\eta_{i}}\neq 0$ and $\forall j\neq i$, $\bar{a}_{j\eta_{i}}=0$.
    • If $\eta_{i}$ is in $B$: $\bar{c}_{i\eta_{i}}\neq 0$ and $\forall j\neq i$, $\bar{c}_{j\eta_{i}}=0$.
@@
The rows of $\mathcal{X}$, $\mathcal{V}$ and $\mathcal{W}$ represent elements in $\mathcal{O}$ as remarked before. The matrix $\mathcal{U}$ represents elements in $\mathcal{O}/p\mathcal{O}$. Statement (i) and (ii) prove that $\overline{\mathcal{V}}$ and $\mathcal{U}$ are linearly independent. This, together with (iv) and (v) guarantee that the collection as in (4) is a basis for $I_{p}/pI_{p}$. The tuples $\gamma$ and $\beta$ give the coordinates of a witness element $\beta_{W}$ in $I_{p}/pI_{p}$ with respect to this basis. Statement (vi) together with (iii) imply that $\{\varphi(\bar{x}_{1})(\beta_{W}),\ldots,\varphi(\bar{x}_{m})(\beta_{W}))\}$ is linearly independent, establishing the $p$-maximality of $\mathcal{O}$.
@@
(p : ℕ) [hpI : Fact \$ Nat.Prime p] ($\mathcal{O}$ : Subalgebra K) ($\mathcal{O}$m : Subalgebra K) (hm : $\mathcal{O}$ $\mathcal{O}$m) where
It includes as part of its fields a term of type TimesTable (Fin (m + n)) $\mathcal{O}$, a term of type Basis (Fin (m + n)) $\mathcal{O}$m (which we do not know explicitly, of course, but serves as a proof that, as a $\mathbf{Z}$-module, $\mathcal{O}_{m}$ is free, finite and has rank $r$), and the certifying data described earlier, together with proofs of the verification statements.
The statements (iv), (v), and (vi), which involve $r$ identities in $\mathcal{O}/p\mathcal{O}$ and $r$ in $\mathcal{O}$, are stated in terms of lists over $\mathbf{F}_{p}$ and $\mathbf{Z}$, respectively, as described in Section 4.2. The proofs can then be solved using decide and/or a repeated use of rfl. From a term of this type, we get a proof of piMaximal p $\mathcal{O}$.
In case a witness $\beta_{W}$ does not exist or cannot be found, we formalized a longer certification scheme for which a certificate always exists when $K$ is a number field, $\mathcal{O}_{m}=\mathcal{O}_{K}$, and $\mathcal{O}$ is $p$-maximal. This requires verifying $r^{2}$ identities in $\mathcal{O}$. For details, we refer to the Appendix A. Additionally, in cases where $m=0$ (corresponding to $p$ being unramified in $K$) a simplified certificate is available.

### v2

@@ 5.6. The Non-monogenic Case
In this section, we introduce a certificate for local maximality which can be used even if the Dedekind criterion does not apply. We specialize to $R=\mathbf{Z}$, our main interest. Our formalization efforts were primarily focused on this case, with $\mathbf{F}_{p}$ modeled as ZMod p. However, we formalized many theorems about semilinear maps (introduced in Lean in (Dupuis et al., 2022)) and their kernels –which we use extensively– in a more general setting, laying the groundwork for future extensions.
The strategy used in the Round-2 algorithm is to reduce the computation of $I_{p}$ and $r(I_{p})$ to a computation in linear algebra over the field $\mathbf{F}_{p}$. We will use an analogous strategy to certify $p$-maximality by verifying that the kernel of certain linear map is trivial. Let $\mathcal{O}$ be a commutative ring which is free as a $\mathbf{Z}$-module and of finite rank $r$, and let $p$ be a prime number. The quotient ring $\mathcal{O}/p\mathcal{O}$ is a finite dimensional $\mathbf{F}_{p}$-algebra. Similarly, the quotient $I_{p}/pI_{p}$ is a finite dimensional $\mathbf{F}_{p}$-module (albeit not an $\mathbf{F}_{p}$- algebra). Consider the map from $\mathcal{O}/p\mathcal{O}$ to the $\mathbf{F}_{p}$-linear endomorphisms of $I_{p}/pI_{p}$.
@@
Let $K$ be $\mathbf{Z}$-torsion-free ring containing $\mathcal{O}$. If the kernel of the map $\varphi$ in (3) is trivial, then $r(I_{p})=\mathcal{O}$.
When $K$ is a $\mathbf{Q}$-algebra, this is an if-and-only-if statement. If $\mathcal{O}^{\prime}$ is a subring of $K$ containing $\mathcal{O}$ and of equal rank, then the previous proposition together with Theorem (5.1) ensures that $\mathcal{O}$ is $p$-maximal if the kernel of $\varphi$ is trivial. Therefore, finding a way to certify that $\varphi$ has trivial kernel becomes our main goal.
@@
Thus, $I_{p}/pI_{p}$ (taking $M=I_{p}$ and $R=\mathbf{Z}$) and $O/pO$ (taking $M=\mathcal{O}$ and $R=\mathcal{O}$) are both special cases of this construction, allowing Lean to infer an $\mathbf{F}_{p}$-module instance for each. Furthermore, Lean will automatically synthesize a CommRing instance for $\mathcal{O}/p\mathcal{O}$, which we then use to give it a $\mathbf{F}_{p}$-algebra structure. Another advantage of this unified approach is that we can define associated objects –such as a $\mathbf{Z}/n\mathbf{Z}$-basis for $M/nM$ given a $\mathbf{Z}$-basis for $M$– and apply it to both cases. We also provide DistribMulAction and SMulCommClass instances that act as a shortcut for typeclass inference, and thereby prevent a timeout in downstream code.
To define map (3) in Lean, we used multiple steps. First, for a parameter : $\mathcal{O}$), we defined the map $I_{p}\to I_{p}/pI_{p}$ sending $\beta\mapsto\overline{\alpha\beta}$. Using Quotient.lift we lifted this to a map $I_{p}/pI_{p}\to I_{p}/pI_{p}$. This was bundled with its proofs of linearity into a linear function, giving an element in $\operatorname{End}(I_{p}/pI_{p})$. After defining the corresponding map $\mathcal{O}\to\operatorname{End}(I_{p}/pI_{p})$, lifting again, and proving linearity, we obtain:
@@
Let us return to the matter of certifying that the kernel of $\varphi$ is trivial. Given a $\mathbf{Z}$-basis for an abelian additive group $M$, we obtain a $\mathbf{Z}/n\mathbf{Z}$-basis for $M/nM$ by mapping it through the $\sigma$-semilinear map $M\to M/nM$, with $\sigma:\mathbf{Z}\to\mathbf{Z}/n\mathbf{Z}$. With these bases, the image of an element in $M$ represented by $(x_{1},\ldots,x_{r})\in\mathbf{Z}^{r}$ will have coordinates $(\bar{x}_{1},\ldots,\bar{x}_{r})\in(\mathbf{Z}/n\mathbf{Z})^{r}$ in $M/nM$. In this way, we obtain an $\mathbf{F}_{p}$-basis for $\mathcal{O}/p\mathcal{O}$ from a $\mathbf{Z}$-basis for $\mathcal{O}$. Finding an $\mathbf{F}_{p}$-basis for $I_{p}/pI_{p}$ is more complex. We will make use of the Frobenius endomorphism $\operatorname{Frob}_{p}$, which, in a ring of characteristic $p$, is the $\mathbf{F}_{p}$-linear map $x\mapsto x^{p}$. For $t\in\mathbf{N}$, the iterated Frobenius endomorphism $\operatorname{Frob}_{p}^{t}$ sends $x\mapsto x^{p^{t}}$. The image of $I_{p}$ under the quotient map $\mathcal{O}\to\mathcal{O}/p\mathcal{O}$ is $I_{p}/p\mathcal{O}$, an $\mathbf{F}_{p}$-subspace of $\mathcal{O}/p\mathcal{O}$. The following proposition lets us realize $I_{p}/p\mathcal{O}$ as the kernel of a linear map:
@@
Let $t$ be an integer such that $r\leq p^{t}$. The kernel of $\operatorname{Frob}_{p}^{t}:\mathcal{O}/p\mathcal{O}\to\mathcal{O}/p\mathcal{O}$ is equal to $I_{p}/p\mathcal{O}$.
While we could theoretically find a matrix representation for this map, an algorithm for computing the kernel of a matrix has not yet, as far as we know, been implemented in Lean. However, we can certify that a given set of elements in $\mathcal{O}/p\mathcal{O}$ is a basis for the kernel of $\operatorname{Frob}_{p}^{t}$ as follows:
Denote by $\bar{x}$ the image in $\mathcal{O}/p\mathcal{O}$ of $x$ in $\mathcal{O}$. Provide multisets $\mathcal{V}=\{v_{1},\ldots,v_{m}\}$ and $\mathcal{W}=\{w_{1},\ldots,w_{n}\}$ with elements in $\mathcal{O}$ –where $r=m+n$– such that $\overline{\mathcal{V}}=\{\bar{v}_{1},\ldots,\bar{v}_{m}\}$ and $\mathcal{U}=\{\operatorname{Frob}_{p}^{t}(\bar{w}_{1}),\ldots,\operatorname{Frob}_{p}^{t}(\bar{w}_{n})\}$ are $\mathbf{F}_{p}$-linearly independent and $\operatorname{Frob}_{p}^{t}(\bar{v}_{i})=0$ for all $i$. By a dimension argument, this guarantees that $\{\bar{v}_{1},\ldots,\bar{v}_{m}\}$ is an $\mathbf{F}_{p}$-basis for the kernel of $\operatorname{Frob}_{p}^{t}$ – and thus for $I_{p}/p\mathcal{O}$.
We can represent $\overline{\mathcal{V}}$ as an $m\times r$ matrix over $\mathbf{F}_{p}$, the rows being the coordinates of the elements in $\overline{\mathcal{V}}$ with respect to the basis for $\mathcal{O}/p\mathcal{O}$. Since we are working over a field, we can choose the elements in $\mathcal{V}$ so that the corresponding matrix is in reduced row echelon form. The linear independence of $\overline{\mathcal{V}}$ can then be read off directly, without any determinant calculation (which is rather slow using the current Lean implementation). Similarly, the right choice of $\mathcal{W}$ will make the linear independence of $\mathcal{U}$ immediate to determine.
@@
With a basis for $I_{p}/pI_{p}$, we can now represent an element in this space as an $r$-tuple, and an element in $\operatorname{End}(I_{p}/p_{p})$ as an $r\times r$ matrix over $\mathbf{F}_{p}$. To certify that the kernel of the map $\varphi$ in (3) is trivial, we can provide a multiset $\mathcal{X}=\{x_{1},\ldots,x_{r}\}$ of elements in $\mathcal{O}$ such that $\varphi(\overline{\mathcal{X}})=\{\varphi(\bar{x_{1}}),\ldots,\varphi(\bar{x_{r}})\}$ is linearly independent. This last verification can be simplified by observing that, generically, we expect to find an element $\beta_{W}\in I_{p}/pI_{p}$ such that $\{\varphi(\bar{x}_{1})(\beta_{W}),\ldots,\varphi(\bar{x}_{r})(\beta_{W})\}$ is linearly independent. This $\beta_{W}$ acts as a witness of the linear independence of $\varphi(\overline{\mathcal{X}})$. Furthermore, we can choose $\mathcal{X}$ in a way that makes the linear independence of the multiset easy to verify. This leads us to the following certificate for $p$-maximality.
@@
Consider $\mathcal{O}$ and $\mathcal{O}^{\prime}$ subalgebras of a $\mathbf{Z}$-torsion-free commutative ring such that $\mathcal{O}\subseteq\mathcal{O}^{\prime}$, and both of rank $r$. Let =$\mathcal{B}=\{b_{1},\ldots,b_{r}\}$ be a $\mathbf{Z}$-basis for $\mathcal{O}$. Let $m,n$ and $t$ be non-negative integers with $r=m+n$ and $r\leq p^{t}$. A certificate for $p$-maximality consists of the following data: $m\times r$ matrix $\mathcal{V}$ over $\mathbf{Z}$ $n\times r$ matrix $\mathcal{W}$ over $\mathbf{Z}$ $n\times r$ matrix $\mathcal{U}$ over $\mathbf{F}_{p}$ $m$-tuple $\nu$ over $\mathbf{Z}_{>0}$ $n$-tuple $\omega$ over $\mathbf{Z}_{>0}$ $r\times r$ matrix $\mathcal{X}$ over $\mathbf{Z}$ $m$-tuple $\beta$ over $\mathbf{Z}$ $n$-tuple $\gamma$ over $\mathbf{Z}$ $r\times m$ matrix $\mathbf{a}$ over $\mathbf{Z}$ $r\times n$ matrix $\mathbf{c}$ over $\mathbf{Z}$
@@
    • if $\eta_{i}$ is in $A$: $\overline{a_{i\eta_{i}}}\neq 0$ and $\forall j\neq i$, $\overline{a_{j\eta_{i}}}=0$;
    • if $\eta_{i}$ is in $B$: $\overline{c_{i\eta_{i}}}\neq 0$ and $\forall j\neq i$, $\overline{c_{j\eta_{i}}}=0$.
@@
The rows of $\mathcal{X}$, $\mathcal{V}$ and $\mathcal{W}$ represent elements in $\mathcal{O}$ as remarked before. The matrix $\mathcal{U}$ represents elements in $\mathcal{O}/p\mathcal{O}$. Statement (i) and (ii) prove that $\overline{\mathcal{V}}$ and $\mathcal{U}$ are linearly independent. This, together with (iv) and (v) guarantee that the collection as in (4) is a basis for $I_{p}/pI_{p}$. The tuples $\gamma$ and $\beta$ give the coordinates of a witness element $\beta_{W}$ in $I_{p}/pI_{p}$ with respect to this basis. Statement (vi) together with (iii) imply that $\{\varphi(\bar{x}_{1})(\beta_{W}),\ldots,\varphi(\bar{x}_{r})(\beta_{W})\}$ is linearly independent, establishing the $p$-maximality of $\mathcal{O}$.
@@
(p : ℕ) [hpI : Fact \$ Nat.Prime p] ($\mathcal{O}$ : Subalgebra K) ($\mathcal{O}$ : Subalgebra K) (hm : $\mathcal{O}$ $\mathcal{O}$’) where
It includes as part of its fields a term of type TimesTable (Fin (m + n)) $\mathcal{O}$, a term of type Basis (Fin (m + n)) $\mathcal{O}^{\prime}$ (which we do not know explicitly, of course, but serves as a proof that, as a $\mathbf{Z}$-module, $\mathcal{O}^{\prime}$ is free, finite and has rank $r$), and the certifying data described earlier, together with proofs of the verification statements.
The statements (iv), (v), and (vi), which involve $r$ identities in $\mathcal{O}/p\mathcal{O}$ and $r$ in $\mathcal{O}$, are stated in terms of lists over $\mathbf{F}_{p}$ and $\mathbf{Z}$, respectively, as described in Section 5.2. The proofs can then be solved using decide and/or decide!. From a term of this type, we get a proof of piMaximal p $\mathcal{O}$.
In case a witness $\beta_{W}$ does not exist or cannot be found, we formalized a longer certification scheme for which a certificate always exists when $K$ is a number field, $\mathcal{O}^{\prime}=\mathcal{O}_{K}$, and $\mathcal{O}$ is $p$-maximal. This requires verifying $r^{2}$ identities in $\mathcal{O}$. For details, we refer to Appendix A. Additionally, in cases where $m=0$ (corresponding to $p$ being unramified in $K$) a simplified certificate is available.